OpenStudy (anonymous):

how do you write -x+2y=4 in intercept form?

OpenStudy (skullpatrol):

Any ideas? :-)

OpenStudy (anonymous):

-x + 2y = 4 +x +x 2y = x + 4 /2 /2 y = x + 2 <------

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

yes!